Kate and Jean paint a house together in 6 days. If they work alone, Kate takes 5 fewer days than Jean would take to do the same job. Which equation represents the time it takes Kate to paint the house by herself?
1) X^2-17x+30= 0
2) x^2-5x-6= 0
3) 2x-1= 0
4) (X^2-5x)/6= 0

Answer by richwmiller(17219)

1/x+1/(x+5)=1/6
(6*x*(x+5))/x+(6*x*(x+5))/(x+5)=(6*x*(x+5))/6
(6*(x+5))+6*x=x*(x+5)
6*x+30+6*x=x^2+5*x
12*x+30=x^2+5*x
x^2-7*x-30=0
(x-10)*(x+3)
(10,-3)
10 days for kate and 15 days for jean
You must have copied incorrectly
The closest of your choices is 1)